OpenStudy (anonymous):

The tangent to the curve y=(sinx)/(1+tanx) at x =0 is parallel to y axis parallel to x axis

OpenStudy (anonymous):

@satellite73 @FoolForMath

OpenStudy (anonymous):

take derivative, replace \(x\) by 0

OpenStudy (anonymous):

yes i got the answer as 0 now wat to do

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

are you sure it is zero? hold on a sec

OpenStudy (anonymous):

u plzz check it

OpenStudy (anonymous):

i think the slope is 1, so it is neither parallel to y or x axis

OpenStudy (anonymous):

if the answer is one then it is equally inclined to x and y axis...

OpenStudy (anonymous):

??? can u tell me

OpenStudy (anonymous):

@experimentX

OpenStudy (anonymous):

@eseidl

OpenStudy (anonymous):

if the answer is one then it is equally inclined to x and y axis...???????

OpenStudy (anonymous):

The derivative is: \[y'=\frac{(cosx)(1+tanx)+sinx(\sec^2x)}{(1+tanx^2)}\]at x=0 we get\[f'(0)=\frac{1}{1}=1\]So, yeah, it is neither parallel to the x-axis nor parallel to the y-axis as satellite73 stated.

OpenStudy (anonymous):

so is it equally inclined to x and y axis

OpenStudy (anonymous):

yeah. I also have a typo in my answer. the denominator should be (1+tanx)^2

OpenStudy (anonymous):

o is it equally inclined to x and y axis

OpenStudy (anonymous):

The slope of the tangent is parallel to the line y=x. You could also say it is equally inclined to the x and y axes.

OpenStudy (anonymous):

if it is 0 then slope is parallel to x axis??????

OpenStudy (anonymous):

If the slope of the tangent was equal to zero, then yes, it would be parallel to the x-axis

OpenStudy (anonymous):

In this case the slope is parallel to the line y=x (actually it is parallel to any line y=x+b, b=constant).

OpenStudy (anonymous):

so it wont be equally inclined to x and y axis? so the answer is none of these

OpenStudy (anonymous):

plzz add this option too equally inclined to x and y axis

OpenStudy (anonymous):

As I said, it is equally inclined to the x and y-axes.

OpenStudy (experimentx):